OpenStudy (anonymous):

Evaluate 32 + (7 − 2) ⋅ 4 − 6 over 3. 24 54 27 28

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (chrisplusian):

Order of operations... what do you start with?

OpenStudy (chrisplusian):

Parentheses so.... 7-2 =5

OpenStudy (anonymous):

32+5

OpenStudy (chrisplusian):

Then you have 32+5*6*4-6 over 3

OpenStudy (chrisplusian):

You can look at the numerator as being in parentheses and then divided by the bottom

OpenStudy (anonymous):

37 24

OpenStudy (chrisplusian):

Next you have to follow from left to right in the numerator... so you do multiplication first. Multiplying 5*6 you get 30

OpenStudy (anonymous):

30*4=120

OpenStudy (chrisplusian):

Then multiplying 30 by 4 you get 120

OpenStudy (anonymous):

2

OpenStudy (chrisplusian):

Then you do addition left to right

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

i suck at math really bad

OpenStudy (chrisplusian):

Ok for clarification....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what number would we do left to right?

OpenStudy (chrisplusian):

\[\frac{ 32+(7-2)*4-6 }{ 3 }\]

OpenStudy (chrisplusian):

Is that your problem?

OpenStudy (anonymous):

yes

OpenStudy (chrisplusian):

I made a mistake when I started cause I added something that didn;t belong there I aplologize

OpenStudy (anonymous):

thats fine :)

OpenStudy (chrisplusian):

But if the problem I just posted is correct you don't have the answer listed because it would be a decimal. Can you check to be sure that what I posted is actually your problem?

OpenStudy (anonymous):

ok heres my problem

OpenStudy (anonymous):